# Memory-Enhanced Neural Solvers for Efficient Adaptation in Combinatorial Optimization

### v2

@@ Appendix G Ablation Study: MEMENTO input features
In Section 3.2, we present MEMENTO, in particular, we present all the inputs that are used by the neural module to derive the new action logits from the memory data. These inputs, or features, are information associated to each past action taken, and that help decide whether those actions should be taken again or not. In Section 4.1, we compare the rule learned by MEMENTO compared to the policy-gradient update REINFORCE. This comparison is made on the features that both REINFORCE and MEMENTO use: i.e. the action log probability, and the return (or advantage if a baseline is used). Although REINFORCE only uses those two features, MEMENTO uses more, which enables to refine even more the update, and also to adapt it to the budget remaining in the search process. As a recall, in addition to the log probability and return, MEMENTO uses: the log probability of the full trajectory, the log probability of the rest of the trajectory after the action was taken, the budget at the time the action was taken, the action logit suggested by MEMENTO when the action was taken, and the budget currently remaining.
To validate the impact of all the features used in the memory, we provide an ablation study of those features. To highlight the interest of using all the additional features, we report the performance of MEMENTO using only the return and the log probability, against the performance of MEMENTO with all features, on Fig. 8. We also report on Fig. 9 a bigger ablation where components are added one after the others.
We can extract two main observations: (i) first, adding the remaining budget completely changes the strategy. We can see on the right panel of Fig. 8 that having access to this additional feature enables MEMENTO to explore much more, and then to focus on high-performing solutions when it gets closer to the end of the budget; (ii) then, Fig. 9 confirms empirically that all features contribute to improving the overall performance.
@@
Figure 9: Full ablation study of MEMENTO: comparing the impact of memory features that are not available in usual policy gradient estimations methods. The left plot reports the best solution found so far. The right plot shows the performance of the latest solution sampled. A: return + logp; B: A + remaining budget; C: B + budget when action was taken; D: C + memory logp; E: D + trajectory logp; F: E + end trajectory logp.

### v3

@@ Appendix G Ablation Study: MEMENTO input features
In Section˜3.2, we present memento, in particular, we present all the inputs that are used by the neural module to derive the new action logits from the memory data. These inputs, or features, are information associated to each past action taken, and that help decide whether those actions should be taken again or not. In Section˜4.1, we compare the rule learned by memento compared to the policy-gradient update REINFORCE. This comparison is made on the features that both REINFORCE and memento use: i.e. the action log probability, and the return (or advantage if a baseline is used). Although REINFORCE only uses those two features, memento uses more, which enables to refine even more the update, and also to adapt it to the budget remaining in the search process. As a recall, in addition to the log probability and return, memento uses: the log probability of the full trajectory, the log probability of the rest of the trajectory after the action was taken, the budget at the time the action was taken, the action logit suggested by memento when the action was taken, and the budget currently remaining.
To validate the impact of all the features used in the memory, we provide an ablation study of those features. To highlight the interest of using all the additional features, we report the performance of memento using only the return and the log probability, against the performance of memento with all features, on Fig.˜8. We also report on Fig.˜9 a bigger ablation where components are added one after the others.
We can extract two main observations: (i) first, adding the remaining budget completely changes the strategy. We can see on the right panel of Fig.˜8 that having access to this additional feature enables memento to explore much more, and then to focus on high-performing solutions when it gets closer to the end of the budget; (ii) then, Fig.˜9 confirms empirically that all features contribute to improving the overall performance.
@@
Figure 9: Full ablation study of memento: comparing the impact of memory features that are not available in usual policy gradient estimations methods. The left plot reports the best solution found so far. The right plot shows the performance of the latest solution sampled. A: return + logp; B: A + remaining budget; C: B + budget when action was taken; D: C + memory logp + trajectory logp + end trajectory logp.
@@ Appendix H Evaluation metrics during training phase
Appendix I Memory retrieval mechanism
In this section, we provide additional motivation for the choice of the node retrieval mechanism used in memento.
Ideally, the retrieval mechanism should retrieve data based on similarity to the current partial solution. But this comes with a computation cost since it requires getting a similarity measure and extracting the k most similar points in the entire memory. We observe that retrieving from the same node is an excellent proxy for similarity, and that the most similar points are very likely to come from the same node. This retrieval strategy hence provides a better trade-off between quality and computation cost, which is why it was selected as the final strategy for memento.
Nevertheless, memento comes as a framework, and each practitioner is free to change part of the method to fit each specific problem. One can hence update the retrieval mechanism if desired. Below is an example of an alternative strategy for retrieval that does consider the partial solution constructed. Since the output of the multi-head attention of pomo ’s decoder builds a low-dimensional representation of the partial solution, one can store this vector in the memory, and when building a new solution, retrieve only the k-nearest neighbors of the current partial solution’s representation. One could even apply further dimensionality reduction to reduce the cost of the nearest neighbor search. This approach brings a significant cost increase, even when using state-of-the-art approximated nearest neighbor JAX implementation. This effect gets worse when batching the attempts, or the problem instances. In the problems and settings considered in this paper, our simplified retrieval approach maintains similar results, while significantly improving scaling.
Appendix J Evolution of MEMENTO learned update over budget
We further analyse how the update rule learned by the MLP evolves over the course of the budget. Specifically, we compute the MLP’s logit correction values across a grid of inputs by varying the normalised return and the log probability of the action, while fixing all the other inputs. These allows us to isolate the learned update pattern across the return and log probability space. As shown in Fig.˜11, memento strongly upweights low-probability, high-return actions, consistent with advantage-weighted updates, but exhibits a sharper concentration around high-value transitions. Importantly, we observe a temporal shift: early in the budget, the MLP assigns nonzero corrections even to uncertain or suboptimal transitions promoting broad exploration. As the remaining budget decreases, the corrections become increasingly selective, focusing on high-return, high-confidence actions.
Figure 11: The update rule discovered by MEMENTO across different budget stages (low, mid, high). The learned update rule shifts from broad exploration to sharper exploitation as the budget decreases.
Appendix K Impact of train-time budget
The main experiments report results for memento trained with a budget of 200 sequential attempts. In this analysis, we evaluate variants of memento trained under alternative training budgets, and evaluate their performance for increasing inference budgets. We report the results in Table 13.
Despite being trained with different budgets, all variants of memento exhibit stable performance across a wide range of inference budgets. This confirms that conditioning on the remaining budget enables flexible behaviour even when the actual inference budget differs from that used in training. Overall, we observe that memento performs robustly across budget mismatches and that all variants generalise reasonably well beyond their training budget. Notably, the model trained with 200 attempts consistently performs the best which we attribute to a tuning bias: the hyperparameters were selected based on validation performance at budget 200, and were not re-tuned for other settings.
Table 13: Performance of memento on TSP100 across different inference budgets, trained with three budget settings.
Inference Budgets
$B=50$ $B=100$ $B=200$ $B=300$ $B=400$ $B=500$ $B=600$ $B=1200$ $B=1600$
Budget Obj. Gap Obj. Gap Obj. Gap Obj. Gap Obj. Gap Obj. Gap Obj. Gap Obj. Gap Obj. Gap
100 7.769 0.0077% 7.767 0.0063% 7.767 0.0058% 7.767 0.0058% 7.767 0.0054% 7.767 0.0056% 7.767 0.0055% 7.767 0.005% 7.766 0.005%
200 7.768 0.0074% 7.767 0.0058% 7.766 0.005% 7.766 0.0042% 7.766 0.0043% 7.766 0.0042% 7.766 0.004% 7.765 0.0035% 7.765 0.0035%
300 7.768 0.0075% 7.768 0.0066% 7.767 0.0056% 7.767 0.0057% 7.767 0.0057% 7.767 0.005% 7.766 0.0043% 7.766 0.0045% 7.766 0.0045%